### v1

# From Fiber to Fabric: Designing the Mechanics of Machine Knitting

Abstract
Knit fabrics are mechanically durable and tough while sufficiently flexible to conform to curved substrates such as the human body. Recent advancements in industrial knitting enable unprecedented control over the pattern design and functionality of next generation knit fabrics. However, the ability to leverage this granular control to predict and tune the mechanical behavior of these fabrics remains limited due to their complex hierarchical and entangled microstructure. This study establishes a comprehensive experimental and numerical framework to characterize and model the mechanical properties of industrial knitted fabrics. By integrating precise experiments, finite element modeling, and a strain energy-based homogenization approach, we provide an accurate demonstration of how stitch length, pattern, and yarn material govern the anisotropic mechanical response of knitted fabrics. In doing so, we establish clear quantitative links between these parameters and key material properties, including stiffness and anisotropy, that emerge directly from our energy-based model. Recognizing that industrial knitting systems can produce fabrics with spatially varying parameters, we extend our framework to predict heterogeneous knits. We show that material transitions have minimal impact on the fabric’s overall mechanical response, so heterogeneous fabrics can be modeled with our framework as patchworks of homogeneous samples. We design the first industrially knit sleeve explicitly optimized for both fit and function, using heterogeneous patterns that conform to target arm musculature and ensure uniform stress distribution. This work bridges the gap between computational modeling and scalable manufacturing, unlocking new possibilities for wearable devices, assistive textiles, and functional applications.
Introduction
Knit fabrics have been used for millennia, valued for their mechanical robustness, ease of manufacturing, and expansive design space. As with other highly entangled material architectures, a knit’s appearance, texture, and mechanical properties are predominately govern by its microstructure [Liu2019, moestopo2020pushing, wang2021structured, surjadi2024double, zhou2024polycatenated]. Knitting is essentially an additive manufacturing process: yarns, spun from fibers, are interlocked in periodic patterns to form the fabric [Singal2024]. The friction and sliding between stitches allow knits to endure repeated large stretches, multiple washing and drying cycles, and harsh environmental conditions while maintaining durability over decades of use without the need for adhesives or additives [warren2018]. While knitting as a technique to produce textiles through the interlace of yarn loops have remained largely unchanged, industrialization has revolutionized the speed of production and the complexity of the resulting textiles in terms of material, heterogeneity and stitch density [sanchez20233d, Collier1999].
Industrial knitting machines and computer-aided design tools have enabled unprecedented control and complexity of the manufacturing process and of the design space [Tessmer2019]. This in turn highlights the potential for research and innovation of modern knits beyond the fashion industry [Popescu2019]. Unlike other fabrication methods, knitting offers a unique synergy between mechanics research and scalable manufacturing [Popescu2019]. The same equipment used for development is utilized in production, reducing the need for process translation and enabling quick transitions from concept to final product. Designers have the ability to not only program the movement and action of every stitch, but also precisely control the timing of these actions. However, the advanced capabilities of industrial knitting are often underutilized due to our limited understanding of the mechanics of knits.
Accurate modeling of the mechanics of knits is challenging because their complex microstructure is based on the many interlocking and sliding elements (Fig. 1A) [mahadevan2024knittingmultistability, Ding2024, wadekar2020optimized, sperl2022estimation, dinh2018prediction, hong2002theoretical]. This is true for even the simplest homogeneous knits, and significantly more so with real-world fabrics that may be constructed from multiple layers of knits with heterogeneous patterns [Ding2024]. The two primary approaches for computational
knit modeling aim to represent either their qualitative or quantitative behaviors. The qualitative approach emphasizes capturing the accurate kinematics and animation of knit fabric such as draping and clothing through homogenization [Singal2024, Sperl2020, Liu2019, Kaldor2008, Kaldor2010, Cirio2017], using tools such as CLO3D [Huang2022]. The quantitative approach aims to capture the physical mechanics at the different knit hierarchies from clasps with deformable silicone yarns [McKee2017, Grandgeorge2021, Baek2021] to loose knits with isotropic yarns, such as monofilament synthetic yarn and shape memory alloy [Poincloux2018, abel2012two]. Multiscale modeling using homogenization can also be applied for physics-based methods [weeger2018nonlinear, dinh2018prediction, fillep2017towards].
In this work, we leverage advances in both approaches and demonstrate a mechanistic understanding of industrial knit fabrics through numerical modeling and precision experiments. We begin by homogenizing the anisotropic mechanical behavior of the different types of yarns. We then incorporate the yarn mechanics in a finite element-based (FE) volumetric model of representative knit stitches (Fig. 1A). We experimentally validate this model through a parametric study, where we explore isolated knit variations by fabricating, pre-processing, and testing homogeneous fabrics with different stitch lengths, knit patterns, and yarn materials (Fig. 1B). Armed with both experimental and numerical results, we abstract the behavior of homogeneous fabrics using a simplified strain energy model that can be used in commercial FE tools to define the relationship between knit parameters and mechanical properties. Having quantified the effects of knit variations, we extend our framework to capture the mechanical response of knit transitions between variations. Finally, we design a heterogeneous knit sleeve that adapts to significant shape changes along the arm to provide constant pressure (Fig. 1C). Building on previous work, we employ homogenization methods to abstract microscopic behavior, enabling an efficient approach that can model both simple, homogeneous swatches and complex, real-world heterogeneous swatches, and ultimately facilitate the modeling and optimization of advanced functional garments.
Mechanics of Yarns.
Before modeling the behavior of knits, we start by characterizing the mechanical behavior of different types of yarns. We choose three types of yarns that are commonly used with knitting machines. Cotton is a 2-ply spun cellulose yarn with a mercerized finish. It consists of short staple natural fibers. Nylon is a synthetic extruded filament with high stretch. This stretchability arises from the many kinked continuous synthetic fibers running from end to end. Polyethylene terephthalate or PET is a stiff monofilament similar to fishing line.
Observing that the yarns behave differently in longitudinal and radial directions, we characterize this anisotropy experimentally [el2023analysis]. The yarns are tensioned lengthwise, and compressed transversely as shown in Fig. 2. (Experiment setup based on Singal et al. [Singal2024]; see SI Section 2.) Nylon and cotton behave with transverse isotropy, where the cross-section is more compliant by four orders of magnitude than the length direction due to the internal voids created by the fiber bundles (Fig. 2). PET, a monofilament, is isotropic.
In physical knit specimens, the diameter of a yarn in the knit can differ significantly from that of a free-standing yarn due to the tension applied during knitting. Therefore, we measure the diameter of the yarn at rest and the diameter of the same type of yarn once knitted . By comparing these measurements, we estimate the tension in the yarn within the knit fabric. The compression tests are conducted on pre-tensioned yarns. Using this experimental data, we homogenize the behavior of individual fibers and numerically model yarns in a knit as a solid curved rod.
Mechanics of homogeneous knit fabric
A typical industrial knit fabric consists of different homogeneous knit units, each defined by varying knit parameters to provide target mechanical behaviors such as stiffness and stretchability. We begin by characterizing the behavior of these homogeneous knit units. To establish an experimental protocol that can be approximated in numerical modeling, we conduct a series of preliminary tests to identify key factors influencing a knit’s mechanical response. These include: cyclic loading, washing and drying, viscoelasticity, and rupture. The preliminary test results are included in SI Section 5.
In these preliminary tests, we apply an equibiaxial strain of 10% to cotton specimens at a rate of 0.1 mm/s. The stress-strain curve of the fabric shows significant dependence on the maximum displacement previously experienced (SI Fig. S7). Subsequent loading to the same displacement exhibits considerable softening and reduced energy dissipation [mullins1969softening]. After washing and drying, the initial behavior is partially recovered, where subsequent testing shows the same softening behavior. Viscoelasticity does not have a significant effect when the displacement rate is varied from 0.1 to 0.4. Premature rupture was observed at the corners for square specimens. Therefore a cruciform geometry is used in subsequent experiments. The details of the tests can be found in SI Section 3.
We experimentally and numerically investigate the influence of two critical manufacturing parameters in knitting in addition to yarn materials. First, stitch length, dictated by the tension during knitting, directly affects the density of the fabric and its mechanical behavior. We consider three distinct stitch lengths (SL): loose, tight, and very tight (Fig. 1B). Physically, they range between settings of 10, 11, and 12, which are equivalent to stitch arc lengths of 3.89 mm, 4.71 mm, and 5.92 mm in Jersey, respectively. These lengths are machine and pattern dependent. Second, knitting patterns emerge from specific arrangements of knits and purls. Knits are pulled towards the face of the fabric, while purls are pulled towards the back. We examine four basic knit structures: single knit Jersey, Garter, Rib, and Seed (Fig. 1B). A Jersey pattern uses identical, Garter alternates rows, Rib alternates columns, and Seed alternates both. Column alternations are in the course direction, which refers to rows of stitches formed by continuous yarn running horizontally. Row alternations are in the wale direction, which corresponds to columns of stitches aligned vertically (Fig. 1B). The sample preparation and testing procedures are in Materials and Methods.
Numerical modeling
Our FE model captures the mechanical behavior of knits with different stitch lengths, pattern and yarn material. The knit geometry and topology is defined using explicit parametric equation of the yarn centerline [crane2023simple]. Instead of modeling individual fibers, the mechanical behavior of the yarns (discussed above) is homogenized as an isotropic (PET) or transversely isotropic (cotton and nylon) material [Poincloux2018, Singal2024]. Starting with the geometry of Jersey as a trigonometric curve [crane2023simple], we modified the geometries of other patterns to match the topology while ensuring that the curve remains continuous and differentiable. A solid geometry is built from the centerline, and meshed using quadratic tetrahedron elements (C3D10). The geometric parameters are determined from microscopic image of knit.
To render the simulations computationally feasible while realistically accounting for critical features such as yarn cross-sectional deformation, frictional contact, sliding, and stitch topologies, we define a representative unit to exploit a knit’s periodicity (Fig. 3A). A minimum of four knots (22) is required to represent the different stitch patterns. To apply periodic boundary conditions, we identify periodic geometry pairs (2 pairs in course and 4 in wale direction), and match the center nodes and planar orientations within (see the matching colored edges in Fig. 3A).
Macroscopic displacements are applied to both course and wale directions and the reaction forces are output during the load step. To address the difference in yarn diameter between its rest and knit state, a prior step is added to the model, where the yarn diameter is gradually increased using a temperature field to its final size [Smith2014], resulting in a variable cross-section depending on contact.
A custom biaxial stage applies equibiaxial stretch on the cruciform samples (Fig. 3B). An overhead camera (Raspberry Pi Camera Module 2) captures the deformation of the center square of the cruciform.
Strain Energy Model
The mechanical responses of knit fabrics under biaxial stress are nonlinear and anistropic (Fig. 3C). To facilitate analysis and isolate the effect of the knit parameters, we use the anisotropic part of a strain energy model based on the Holzapfel-Gasser-Ogden model [holzapfel2000new], which is already implemented in commercial FE software including Abaqus:
where is an indicator of initial stiffness, is a general indicator of stiffness, and is an indicator of anisotropy (Fig. 3D). As approaches 45∘, the more the behavior in the course and wale directions become similar. is the fourth stretch invariant, defined as:
where is the deformation tensor, is the direction of anisotropy, and , are the stretches in the course and wale directions, respectively. In this model, we assume plane stress with zero stress out-of-plane, so that the effect of anisotropy is planar. The corresponding first Piola-Kirchhoff stress (nominal stress) is therefore:
In the low-stress regime with 1, this strain energy can be approximated by . As illustrated in Eqs. 1 and 3, both the strain energy and stress expressions include terms from the course and wale directions (). This indicates that deformations in the course and wale directions are coupled.
We compare the behavior of fabrics depending on stitch length, pattern, and material using , , and . The parameters are fitted to the stress-strain curves obtained experimentally and through simulations. We use the Mesh Adaptive Direct Search (MADS) algorithm to find the optimal parameters by minimizing the squared-difference of the simplified model and the stress-strain response [nomad4paper]:
where is the sample size, , represents the measured nominal stress for the th sample, and denotes the resulting nominal stress from the strain model with the given parameter .
Results from the Systematic Testing of Knitting Variables
We argue that the mechanical behavior of the knit fabric becomes most representative after preliminary loading, once the fabric has stabilized and further testing yields no significant changes. In this stabilized state, we perform two equibiaxial experiments on fabric samples for each variation in stitch length, pattern, and material. The results are shown as stress-strain curves of the experimental, simulated, and fit data in Fig. 4. The responses are fit using the strain energy model and shown in Fig. 5. All results are shown for the course direction only; results for the wale direction are in SI Section 6.
Figure 5A highlights that stitch length primarily influences overall stiffness (), with shorter stitch lengths increasing stiffness. Altering the pattern has the greatest impact on fabric anisotropy, as changes in pattern directly affect the symmetry of the fabric (Figure 5B). Material variations affect both initial stiffness () and overall stiffness () by an order of magnitude, but not anisotropy (Figure 5C).
These results make clear that each variable — stitch length, pattern, and material — affects different aspects of the mechanical response. Collectively, they offer practical design guidelines for tailoring the behavior of homogeneous fabrics. Specifically, material selection defines the overall stiffness range, pattern choice determines the degree of anisotropy, and stitch length serves as a fine-tuning parameter to achieve the desired stiffness response.
Numerical validation
The numerical results and the proposed strain energy model were validated against experimental data by calculating the normalized root mean squared error (NRMSE), normalized by the maximum stress value, for each variation. The normalized error was below 5%, with two exceptions for all simulations and strain energy models. At a stitch length of 10, the simulation error reached 12.3%, attributed to the tightly knotted configuration leading to increased contact interference within the knit structure. At a stitch length of 12, the model’s fit error was 10.7%, reflecting the model’s compromise in fitting the strain energy across both the course and wale directions. This balancing act can introduce discrepancies, as the model aims to capture behavior across multiple deformation modes.
Heterogeneous Fabrics
Industrial knitting machines can vary the above parameters within the same piece of fabric. In fact, heterogeneous fabrics are necessary to achieve spatially varying functionalities. We begin by studying the effect of spatially transitioning from one parameter to another within the same fabric.
Heterogeneous variations can occur both in course and wale directions (see diagrams in Fig. 6). Change in material in the course direction requires a transition technique called intarsia, where new variation is knit into a new column (see drawing in Fig. 6). In the wale direction, the change occurs seamlessly, since a row of knots of a length, pattern, or material can be directly knit into a different row.
Here, we present results for material changes, which had the largest impact on mechanical behavior of homogeneous specimens. The results for variations in patterns are given in SI Section 6. As we aim to isolate the effect of the transition from one mechanical regime to another, we impose uniaxial stress on the samples. We subject an initially square knit fabric () using two different yarn materials (nylon and cotton), varied in the course and wale directions, up to 100% strain in both course and wale directions (total of four tests). When the sample is tested such that both materials are clamped together on both sides (material change is in line with the testing axis), we say the sample is tested in parallel. Conversely, when the material change is perpendicular to the testing axis, we say the sample is tested in series. For the in parallel test scenario, both fabrics are subjected to the same amount of strain. For the in series test scenario, we use digital image correlation (DIC) to track the displacement of the transition line between materials to extract the strain experienced by the two parts of the fabric.
We compare the resulting behavior against those of homogeneous fabrics under the same testing condition to extract the influence (if any) of the transitions. When testing in parallel (Fig. 6 top row), we see that the heterogeneous response is closely approximated by the average response of homogeneous swatches. When testing in series (Fig. 6 bottom row), we see that the individual material response is in line with the individual testing of homogeneous swatches, showing that material transition does not influence the behavior of the fabric.
Design of a uniform stress sleeve
For skin-tight garments such as bodysuits, compression sleeves, and leggings, some stretch is present throughout the fabric to reduce wrinkling. To account for the varying stretch across different parts of the human body, e.g., different circumferences between the bicep and the wrist, additional panels are often sewn in. These extra panels reduce fabric damage from excess stretch and alleviate discomfort associated with large compressive stress to the body. (As shown earlier, generally, stress increases exponentially with stretch.) However, achieving uniform compression with paneling is challenging and produces fabric waste. The additional seams can also cause skin irritation and introduce weaknesses in the garments [Xiong2018]. Building on the results above, we demonstrate a single piece of simply shaped heterogeneous fabric that can accommodate spatially varying stretches while maintaining the uniform stress distribution throughout. In contrast to previous work that uses thermal actuation for self-fitting after production [Tessmer2019, Granberry2019, Eschen2018], our approach is to optimize the topology a priori for fit and uniform stress distribution.
Using the arm as an example, we reconstruct the arm of a muscular volunteer in three dimensions. By generating a quadrilateral mesh and mapping it onto a regular cylinder with a radius matching that of the wrist, we derive the strain experienced at each material point on the arm (Fig. 7A). We then identify knit parameter combinations that produce consistent stress under these varying strains (Fig. 7B). This results in as uniform stress as possible, even though different parts of the sleeve may experience dissimilar strains when it is worn. We verify that the fabric stretches as predicted by the design on the scanned arm, with no slack observed (Fig. 7D).
This work establishes a robust and accurate experimental and numerical framework for predicting and tuning the mechanical behavior of industrial knit fabrics, enabling precise control over stiffness, anisotropy, and spatially varying functionalities. By integrating computational modeling, simplified energy-based homogenization, and scalable manufacturing, our approach enables the streamlined design and production of optimized, personalized textiles—opening new frontiers in wearable devices, functional fashion, and beyond.
Materials and Methods
Manufacturing of knit specimens
We use a Stoll CMS 330 industrial knitting machine to knit the fabric swatches. The patterns are manually programmed except for the sleeve, which is generated algorithmically. Detailed information is given in SI Section 1.
Yarn experimental testing
The yarns are washed and dried prior to mechanical loading. In tension, the yarns are loaded from the relaxed state (50 mm) until failure at a rate of 0.1 mm/s. The diameter of the yarns are measured throughout. In compression, the yarns are pre-tensioned until their diameter match those of the yarns in knit fabrics. They are then compressed at a rate of 0.01 mm/s up to a threshold of 50 N. A universal testing machine (Instron 68SC) is used in all tests. Additional details on the procedure are given in SI Section 2.
Fabric swatch experimental testing
The homogeneous specimens are laser cut to a cruciform shape such the loads are applied to a 30 30 mm gauge area, except for the PET samples which are cut to a square shape due to unravelling at the corners of the cruciform.
All specimens are washed and dried prior to mechanical testing. Optical microscopy is used to obtain the geometrical parameters for FEM. All homogeneous specimens are tested using a custom biaxial testing stage (Fig. 3B). The specimens are clamped on all four edges. Equibiaxial strain is applied up to 30%. The actual strain in the gauge area is captured with an overhead camera. Three specimens are loaded cyclically three times each for each parameter variation.
Uniaxial testing is conducted to measure the influence of the different transitions (heterogeneous samples) using specimens of 50 50 mm. Further details on the testing procedures can be found in SI Sections 3 and 4. Both are based on previous testing procedures described by Connolly et al. [Connolly2019].
Numerical modeling of knits
ABAQUS/Standard 2022 is used to simulate the mechanical behavior of knit fabrics. The rest configuration of the knit geometry is described by a space curve parameterized along its arc length. A 2 2 representative unit is defined and meshed using quadratic tetrahedral elements. Two static loading steps with nonlinear geometry are defined: The first uses thermal expansion to prevent any interference while allowing proper contact between the yarns. The second applies prescribed displacements in the two directions. This procedure is detailed in SI Section 7.
Design of the sleeve
A 3D scanned arm is used as the target shape. A cylinder with a circumference equal to that of the wrist is used as the initial geometry. The resulting strain on each mesh is estimated by generating a structured quadrilateral mesh on both the initial and target geometries with the same number of elements and a one-to-one mapping. The yarn material is initialized as cotton for all meshes. Due to peak stresses exceeding the yield stress of our cotton yarn, we switched certain regions to Rib nylon, which is more compliant. The homogenized model is used to predict the resulting stress based on the knit parameters–stitch length, pattern, and material–at these estimated strains. The parameters minimize variance in stress distribution throughout the fabric. The optimized parameters are filtered to obtain contiguous parameter regions. The details in selecting excessive strain region, filtering and optimization are state in SI Section 9, fabrication and experiments. The final design is processed using CREATE PLUS (Stoll) and knitted using the Stoll CMS 330.
Acknowledgments
We thank Lavender Tessmer for her guidance and early collaboration in industrial knitting. We thank Ushanth Balasuriya for volunteering his arm. Some of the computing for this project was performed on the Stanford University Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results. We also thank Conor O’Brien at Otherlab for his advice and support in knitting the sleeve.